OpenStudy (anonymous):

2X^3 + 54 . FACTOR COMPLETELY

jimthompson5910 (jim_thompson5910):

2x^3 + 54 2(x^3 + 27) Now use the sum of cubes factoring rule.

OpenStudy (anonymous):

Here is an easier way. 2x^3+54=0. Divide both sides by 2. Now you have x^3+27=0. Subtract 27 from both sides. Now you have x^3=-27. Now you just square root both sides and you get x=-3

jimthompson5910 (jim_thompson5910):

kirk.freedman he wants to factor, not solve

OpenStudy (anonymous):

Ohh, sorry did not see that part.

jimthompson5910 (jim_thompson5910):

no worries